# Large Language Models for Blockchain Security: A Systematic Literature Review

### v4

@@ 1 Introduction
As the digital era advances, the confluence of artificial intelligence with blockchain technology emerges as a groundbreaking development, particularly at the juncture where Large Language Models (LLMs) [1, 2, 3, 4, 5, 6] intersect with the ever-evolving domain of blockchain security [7, 8, 5, 6, 9, 10]. LLMs have risen to the forefront of blockchain security [11, 12], showcasing profound capabilities in text generation and comprehension [13, 14, 6, 15]), especially in source code analysis. These abilities mirror human-like proficiency [16, 17]. This transformative impact is attributable to their expansive datasets, sophisticated architectures, and the deep neural networks that underpin their operational frameworks [1, 18, 17].
@@
Domains | Amounts | Publications | |||||
Smart Contract Auditing | 15 |
Block Transaction Detection | 2 | BLOCKGPT [26] and Nicholls et al. [43] | |||||
Contract Dynamic Analysis | 2 | LLM4FUZZ [44] and ACFIX [45]. | |||||
@@ 2.2.1 Key Components of Blockchain Security
Smart Contract Security: Smart contracts are self-executing contracts with predefined rules and conditions encoded on the blockchain. Ensuring the security of smart contracts is essential to prevent vulnerabilities, exploits, and unauthorized access [92, 93, 94]. Techniques such as formal verification, code auditing, and secure development practices are employed to mitigate risks associated with smart contracts, including reentrancy attacks, integer overflow/underflow, and unchecked external calls [95, 96].
@@ 3.1 LLM as Code auditor on Smart Contracts
The application of LLM in the domain of smart contract code auditing and vulnerability detection can be succinctly encapsulated as follows: Advanced tools, such as SMARTINV [29], GPTScan [13], David et al. [30], Karanjai et al. [31], ContractArmor [32], Ortu et al. [33], ASSBert [34], PSCVFinder [35], LLM4Vuln [36], TrustLLMf [37], AuditGPT [38], Chen et al. [40], Jain et al. [41], PropertyGPT [39] and SolGPT [42]. As shown in Table.2, these tools powered by Large Language Models signify a monumental shift from traditional, pattern-based analysis methodologies towards more contextually aware and comprehensive inspection techniques. These cutting-edge tools extend their analytical prowess beyond static patterns by knitting together disparate threads of information, including the nuanced aspects of natural language documentation that detail the intended functions and transactional constructs of smart contracts.
@@
Research | Functionality | LLM(s) | Publication |
@@
TrustLLMf [37] | Smart contracts audit | CodeLlama-13b | arXiv,2024 |
@@ 3.2 LLM as Analyzers for abnormal transaction
The application of LLMs for blockchain transaction analysis, such as BLOCKGPT [26] and Nicholls et al. [43], underscores their crucial role in conducting real-time monitoring to detect signs of irregular or suspicious behavior. These tools in Table 3 represent a significant advancement in the field, as they provide a more dynamic and adaptable approach to identifying potential threats within blockchain transactions.
@@ 3.3 LLM as Fuzzer for Smart Contract
Large Language Models (LLMs) have been increasingly employed to elevate the process of fuzzing, particularly in the realm of smart contract security analysis, such as LLM4FUZZ [44] and ACFIX [45]. This methodology in Table 4 involves utilizing LLMs to accurately assess the complexity and vulnerability likelihood of specific code regions within a smart contract. Consequently, these metrics serve to guide the direction and focus of fuzzers, steering them toward code segments that are more likely to harbor potential security threats.
@@ 3.4 LLM as Developer for Smart Contract
work | function | LLM(s) | Publication |
@@ 3.5 LLM as Participants for Cryptocurrency community
work | function | LLM(s) | Publication |
@@ 3.6 Miscellaneous
work | function | LLM(s) | Publication |

### v5

@@ Abstract
[inst3]organization=University College London,country=United Kingdom
@@ 1 Introduction
As the digital era advances, the confluence of artificial intelligence with blockchain technology emerges as a groundbreaking development, particularly at the juncture where Large Language Models (LLMs) [1, 2, 3, 4, 5, 6] intersect with the ever-evolving domain of blockchain security [7, 8, 5, 6, 9, 10, 11]. LLMs have risen to the forefront of blockchain security [12, 13], showcasing profound capabilities in text generation and comprehension [14, 15, 6, 16]), especially in source code analysis. These abilities mirror human-like proficiency [17, 18]. This transformative impact is attributable to their expansive datasets, sophisticated architectures, and the deep neural networks that underpin their operational frameworks [1, 19, 18, 20].
@@
| Domains | Amounts | Publications | ||||||||
| Smart Contract Auditing | 22 |
| Block Transaction Detection | 3 | BLOCKGPT [28], Nicholls et al. [53] and ZipZap [54] | ||||||||
| Contract Dynamic Analysis | 3 | LLM4FUZZ [55], ACFIX [56] and Sun et al. [57] | ||||||||
@@ 2.2.1 Key Components of Blockchain Security
Smart Contract Security: Smart contracts are self-executing contracts with predefined rules and conditions encoded on the blockchain. Ensuring the security of smart contracts is essential to prevent vulnerabilities, exploits, and unauthorized access [104, 105, 106, 107]. Techniques such as formal verification, code auditing, and secure development practices are employed to mitigate risks associated with smart contracts, including reentrancy attacks, integer overflow/underflow, and unchecked external calls [108, 109, 52].
@@ 3.1 LLM as Code auditor on Smart Contracts
The application of LLM in the domain of smart contract code auditing and vulnerability detection can be succinctly encapsulated as follows: Advanced tools, such as SMARTINV [31], GPTScan [14], David et al. [32], Karanjai et al. [33], ContractArmor [34], Ortu et al. [35], ASSBert [36], PSCVFinder [37], LLM4Vuln [38], TrustLLM [39], AuditGPT [40], Chen et al. [42], Jain et al. [43], PropertyGPT [41], SolGPT [44], DeFiAligner [45], RepairBench [46], LLM-SmartAudit [47], Hyperion [48], AdvSCanner [49], Sóley [50], Jiang et al.[51] and XPLOGEN [52]. As shown in Table.2, these tools powered by Large Language Models signify a monumental shift from traditional, pattern-based analysis methodologies towards more contextually aware and comprehensive inspection techniques. These cutting-edge tools extend their analytical prowess beyond static patterns by knitting together disparate threads of information, including the nuanced aspects of natural language documentation that detail the intended functions and transactional constructs of smart contracts.
@@
| Research | Functionality | LLM(s) | Publications |
@@
| TrustLLM [39] | Smart contracts audit | CodeLlama-13b | arXiv,2024 |
@@
| DeFiAligner [45] | Logic inconsistency detection | GPT-3.5, GPT-4, GPT-4o | AFT,2024 |
| RepairBench [46] | AI-driven program repair | GPT-4o etc. | arXiv,2024 |
| LLM-SmartAudit [47] | Advanced vulnerability detection | GPT3.5, GPT-4o | arXiv, 2024 |
| Hyperion [48] | DApp inconsistency detection | LLaMA2 | arXiv, 2024 |
| AdvSCanner [49] | Reentrancy vulnerability exploitation | GPT-4o | ASE,2024 |
| Sóley [50] | Logic vulnerability detection | CodeBERT | JSS,2025 |
| Jiang et al.[51] | LLM-based Code Smell Detector | GPT-4 | TSE,2024 |
| XPLOGEN [52] | Exploit Synthesis Tool | GPT-4 | ASE, 2024 |
@@ 3.2 LLM as Analyzers for abnormal transaction
The application of LLMs for blockchain transaction analysis, such as BLOCKGPT [28], Nicholls et al. [53] and ZipZap [54], underscores their crucial role in conducting real-time monitoring to detect signs of irregular or suspicious behavior. These tools in Table 3 represent a significant advancement in the field, as they provide a more dynamic and adaptable approach to identifying potential threats within blockchain transactions.
@@ 3.3 LLM as Fuzzer for Smart Contract
Large Language Models (LLMs) have been increasingly employed to elevate the process of fuzzing, particularly in the realm of smart contract security analysis, such as LLM4FUZZ [55], ACFIX [56] and Sun et al. [57]. This methodology in Table 4 involves utilizing LLMs to accurately assess the complexity and vulnerability likelihood of specific code regions within a smart contract. Consequently, these metrics serve to guide the direction and focus of fuzzers, steering them toward code segments that are more likely to harbor potential security threats.
@@ 3.4 LLM as Developer for Smart Contract
| Research | Functionality | LLM(s) | Publications |
@@ 3.5 LLM as Participants for Cryptocurrency community
| Research | Functionality | LLM(s) | Publications |
@@ 3.6 Miscellaneous
| Research | Functionality | LLM(s) | Publications |